OpenStudy (anonymous):

(1/2a+1/2b)^2 simplify

OpenStudy (king):

4(a^2+b^2+2ab)/4a^2b^2 =(a+b)^2/(ab)^2 =\[(\frac{a+b}{ab})^{2}\]

OpenStudy (anonymous):

what

OpenStudy (king):

(a+b)^2/(ab)^2

OpenStudy (anonymous):

how is that the answer a half+a half squared is equalled to that